# Efficient absorbing boundary conditions for conservation laws: Application to the piston problem of gas dynamics

### v1

@@ 6 Conclusion
Methods from two families of absorbing boundary conditions were studied, and their performance was tested for the piston problem of gas dynamics. The relaxation method with a scalar-valued weight function is the least expensive and simplest to implement. Karni’s slowing-down and damping operators are able to selectively absorb waves from a specific characteristic field, allowing physical waves to circulate back into the computational domain. However, this technique requires the addition of an artificial source term that might become stiff depending on the prescribed damping rate, leading to additional time step size restrictions to maintain stability with an explicit time integrator. Among the extensions for both techniques presented in this work, the relaxation method with a matrix-valued weight function is an attractive alternative, as it preserves the remarkable robustness of the relaxation method (even in the nonlinear setting), while selectively damping outgoing waves. We expect this property to be more advantageous in a case where more interactions take place inside the sponge layer, for instance, when the sponge layer is used as a wavemaker (i.e. nonconstant far-field state), which is a common application for the relaxation method. Furthermore, sophisticated scalar weight functions (e.g. that depend on the conserved quantities), that have already been developed for the relaxation method, can be used with this extension in a straightforward manner. Despite being the least efficient technique considered in this work, we expect the nonlinear damping operator to be more effective for high-order discretizations, where a high-order approximation of the solution is available inside each cell. Thus, more information can be used for the computation of the integral damping term.
@@ Acknowledgements
I would like to express my deepest gratitude to Prof. David Ketcheson and Prof. Giovanni Russo for their invaluable advice and support throughout the development of this work.
Appendix A Appendix
A.1 Construction of damping operators
@@
The construction of the damping operators and the matrix-valued weight function of Sections 2.1, 3.4, and 3.2, relies on the computation of a matrix $\mathbf{M}=\mathbf{R}\mathbf{\Sigma}\mathbf{R}^{-1}$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{i})$, and $\mathbf{R}$ contains the right eigenvectors of the flux Jacobian matrix as columns. For the linear equations (42) we have
@@
A.2 Discretization of integral source term with nonconservative product
@@
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})$ $=\int_{x}^{x_{j}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\sum_{i=j}^{m-1}\int_{x_{i}}^{x_{i+1/2}^{-}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\int_{x_{i+1/2}^{+}}^{x_{i+1}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\int_{\{x_{i+1/2}\}}\mathbf{M}\partial_{z}\mathbf{q}\,dz,$ (57)
[…]
@@
[…]
where $\phi$ is a Lipschitz continuous function connecting $\mathbf{q}^{-}(t)=\lim_{x\to x_{i+1/2}^{-}}\mathbf{q}(x,t)$ with $\mathbf{q}^{+}(t)=\lim_{x\to x_{i+1/2}^{+}}\mathbf{q}(x,t)$ [26]. Here we choose $\phi$ to be simply a straight line $\phi(s)=s\mathbf{q}^{+}+(1-s)\mathbf{q}^{-}$, and approximate the integral in (59) with the midpoint rule.
A.3 Stability of the slowing-down operator for nonlinear equations
One can show that the slowing-down operator does not destroy the $L^{2}$ stability of a linear system. Indeed, we can consider a general linear system of $N$ hyperbolic PDEs
@@ A.3 Stability of the slowing-down operator for nonlinear equations
defined over the real line such that $\lim_{x\to\pm\infty}\mathbf{q}(t,x)=\bf 0$. It is then clear that $\partial_{t}||\mathbf{q}(t,\cdot)||_{L^{2}()}=0$. Let us analyze the effect of the slowing-down operator applied over the sponge layer $[x_{s},x_{\infty}]$. For instance, we absorb the right-going waves, i.e., we solve
@@
where $\mathbf{S}(x)=\mathbf{R}\mathbf{\Lambda}_{\mathbf{S}}(x)\mathbf{R}^{-1}$ such that $\mathbf{A}=\mathbf{R}\mathbf{\Lambda}\mathbf{R}^{-1}$, and the main entries of the diagonal matrix $\mathbf{\Lambda_{S}}(x)$ are given by
$\tilde{\lambda}_{i}(x)=\begin{cases}\lambda_{i},\quad&\text{if }\lambda_{i}\leq 0,\\ s(x)\lambda_{i},\quad&\text{if }\lambda_{i}>0.\end{cases}$ (62)
The scalar smooth slowing-down function $s(x)$ strictly decreases from $1$ to 00 inside the sponge layer $[x_{s},x_{\infty}]$ and is constant everywhere else. Then, multiplying (62) by $\mathbf{q}^{T}$ and integrating in space, we get
[…]
The inequality is a consequence of $\mathbf{S}^{\prime}(x)$ being semi-negative definite, since ${\tilde{\lambda}_{i}}^{\prime}(x)\leq 0$.
@@
for which we know that $\partial_{t}||u||_{L^{2}()}\leq 0$ holds. The modified equation reads
@@
$\frac{1}{2}\partial_{t}\int_{u}^{2}\,dx=-\frac{1}{3}\int_{s}(x)\partial_{x}u^{3}\,dx=-\frac{1}{3}\int_{\partial}{}_{x}[s(x)u^{3}]-s^{\prime}(x)u^{3}\,dx=\frac{1}{3}\int_{x_{s}}^{x_{\infty}}s^{\prime}(x)u^{3}\,dx.$ (66)
If $u<0$ inside the sponge layer, the last term (thus $\partial_{t}||u||_{L^{2}()}$) becomes strictly greater than 00 and the entropy-dissipative nature of the original system is lost.

### v2

@@ 6 Conclusion
Methods from two families of absorbing boundary conditions were studied, and their performance was tested for the piston problem of gas dynamics. The relaxation method with a scalar-valued weight function is the least expensive and simplest to implement. Karni’s slowing-down and damping operators are able to selectively absorb waves from a specific characteristic field, allowing physical waves to circulate back into the computational domain. However, this technique requires the addition of an artificial source term that might become stiff depending on the prescribed damping rate, leading to additional time step size restrictions to maintain stability with an explicit time integrator. Among the extensions for both techniques presented in this work, the relaxation method with a matrix-valued weight function is an attractive alternative, as it preserves the remarkable robustness of the relaxation method (even in the nonlinear setting), while selectively damping outgoing waves. We expect this property to be more advantageous in a case where more interactions take place inside the sponge layer, for instance, when the sponge layer is used as a wavemaker (i.e. nonconstant far-field state), which is a common application for the relaxation method. Furthermore, sophisticated scalar weight functions (e.g. that depend on the conserved quantities), that have already been developed for the relaxation method, can be used with this extension in a straightforward manner. Due to its higher computational cost and lack of robustness for coarse grids, the newly proposed nonlinear damping operator is not competitive with existing ABCs considered in this work. However, this technique could warrant further investigation in scenarios where a higher-order representation of the solution is available inside each cell, since the integral damping term could leverage this additional information. While this work focused on one-dimensional problems, the extension of the matrix-valued relaxation method to multiple spatial dimensions could be achieved for rotationally invariant systems following the approach in [karni_far-field_1996]; this is the subject of future work.
@@ Acknowledgements
I would like to express my deepest gratitude to Prof. David Ketcheson and Prof. Giovanni Russo for their invaluable advice and support throughout the development of this work. This work has been funded by King Abdullah University of Science and Technology.
Appendix A Reflection coefficients
Based on the relations among the different ABCs presented in Section 3, it is possible to provide analytical estimates of the absorbing properties of the SDO and RM techniques in a unified framework. For this, we restrict ourselves to linear systems of the form (2) and begin by considering the SDO 2.1 with no slowing down, i.e., $s_{i}(x)\equiv 1$ for all $i$. From (11) we can see that, if $\int_{x_{s}}^{x}d_{i}(z)\,dz\to+\infty$ as $x\to x_{\infty}$, the right-going information will be completely absorbed inside the sponge layer. In practice, however, $\max(d_{i})=\sigma_{i}$ is chosen to be large but finite to avoid numerical stiffness. Thus, reflection of waves will occur at $x_{\infty}<+\infty$.
In a similar manner as done in [peric2018analytical] by Perić and Abdel-Maksoud for scalar-type damping operators, we assume that the boundary at $x_{\infty}$ is perfectly reflecting to estimate the reflection coefficient of the sponge layer associated with SDO. We consider an initial condition that is compaclty supported in the computational domain $0,x_{s}]$, which has a characteristic wavelength $L$, and is centered at $\mathring{x}$ ($\mathring{x}+L/2<x_{s}$). Moreover, the incoming wave is is purely right-going, i.e.,
${\mathbf{q}}(0,x)=\mathbf{q}^{\text{inc}}(x)=f(x)\chi_{[\mathring{x}-L/2,\mathring{x}+L/2]}(x)\sum_{\lambda_{i}>0}a_{i}\mathbf{r}_{i},$ (58)
for some coefficients $a_{i}$. In characteristic variables, we allow the solution to evolve up to a point where the information in each characteristic field has left the sponge layer, i.e.,
$t_{\infty}=\min_{t>0}\max_{\lambda_{i}>0}(\mathring{x}-L/2+\lambda_{i}t)>x_{\infty}.$
Then, for each $x>x_{\infty}$, and for each characteristic field $i$, we have
$w_{i}(t_{\infty},x)$ $=a_{i}f(x-x_{i})\chi_{[x_{i}-L/2,x_{i}+L/2]}(x)\exp\left(-\int_{x-\lambda_{i}t_{\infty}}^{x}d_{i}(z)\,dz\right)$ (59)
$=a_{i}f(x-x_{i})\chi_{[x_{i}-L/2,x_{i}+L/2]}(x)\exp\left(-\int_{x_{s}}^{x_{\infty}}d_{i}(z)\,dz\right),$ (60)
where $x_{i}=\mathring{x}+\lambda_{i}t_{\infty}$.
Let us assume that the eigenstructure of the system is symmetric, like in the case of the linearized Euler equations in Lagrangian coordinates, presented in Section 4 below. Consequently, the information from each characteristic field $i$ is reflected back through its reciprocal characteristic field $j(i)$ such that $\lambda_{j}=-\lambda_{i}$, after interacting with the perfectly reflecting boundary at $x_{\infty}$. Thus, the total mass (of the $k$-th conserved variable) of the reflected wave is given by
[…]
while the total mass of the incoming wave is
[…]
Then, we define the reflection coefficient associated with the damping operator as
$C_{R}^{\text{SDO}}=\frac{M_{\text{ref}}}{M_{\text{inc}}}=\frac{\sum_{\lambda_{j}<0}|a_{j}||\mathbf{r}_{jk}|\exp\left(-\int_{x_{s}}^{x_{\infty}}d_{i}(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|},$ (65)
which, due to the symmetry of the eigenstructure, can be rewritten as
$C_{R}^{\text{SDO}}=\frac{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|\exp\left(-\int_{x_{s}}^{x_{\infty}}d_{i}(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|}.$ (66)
Following a similar reasoning, we can compute the reflection coefficient of the scalar damping operator (13). In this case, after the wave is reflected at $x_{\infty}$, the left-going information is not only propagated back by left-going characteristics $\lambda_{j}<0$, but also damped at the rate
$\exp\left(-\mathrm{sgn}(\lambda_{j})|\lambda_{j}^{-1}|\int_{x-\lambda_{j}t}^{x}d(z)\,dz\right)=\exp\left(\lambda_{j}^{-1}\int_{x-\lambda_{j}t}^{x}d(z)\,dz\right).$
Hence, the total mass of the reflected wave is
[…]
which, due to the symmetry of the eigenstructure, gives us the reflection coefficient
$C_{R}^{\text{S-SDO}}=\frac{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|\exp\left(-2\lambda_{i}^{-1}\int_{x_{s}}^{x_{\infty}}d(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|}.$ (68)
We can see that the reflection coefficient of S-SDO will be smaller or larger than that of SDO depending on the magnitude of the eigenvalues $\lambda_{i}$.
Using this setup, we can also compute the reflection coefficient for the matrix-valued relaxation method RM-M. We assume that left-going waves are unperturbed ($\Gamma_{j}\equiv 1$ for $j<0$), and consider the damping function (31), which necessarily depends on the discretization through $\Delta t$. Then, proceeding as above, the reflection coefficient of this method is given by
$C_{R}^{\text{RM-M}}$ $=\frac{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|\exp\left(\frac{1}{\Delta t\lambda_{i}}\int_{x_{s}}^{x_{\infty}}\ln\circ\Gamma_{i}(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|}$ (69)
$\approx\frac{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|\exp\left(\frac{x_{\infty}-x_{s}}{C\Delta x}\int_{0}^{1}\ln\circ\Gamma_{i}\circ\phi^{-1}(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|}.$ (70)
Here, we have used the fact that $\phi$ is invertible in $x_{s},x_{\infty}]$ (16), and assumed the CFL condition $C=\Delta t\max_{i,p}|\lambda_{p}|/\Delta x$. The factor $(x_{\infty}-x_{s})/(C\Delta x)$ indicates that the effectiveness of the RM-M method depends not only on the choice of $\Gamma_{i}$, but also on the number of cells in the sponge layer. Similarly as before, we can compute the reflection coefficient for the scalar relaxation method RM, which is given by
$C_{R}^{\text{RM}}\approx\frac{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|\exp\left(2\frac{x_{\infty}-x_{s}}{C\Delta x}\int_{0}^{1}\ln\circ\Gamma_{i}\circ\phi^{-1}(z)\,dz\right)}{\sum_{\lambda_{i}>0}|a_{i}||\mathbf{r}_{ik}|}.$ (71)
As can be expected from our assumptions, if information from all the characteristic fields is damped with the scalar method RM, the reflection coefficients will be smaller than those of the directional methods RM-M.
A.1 Comparsion with numerical results
To validate the theoretical estimates of the reflection coefficients derived in the previous section, we consider the linearized equations (48). For this, we take $x_{s}=20\pi$, different values of $x_{\infty}$. The initial condition is given by the purely right-going pulse
$\mathbf{q}(x,0)=\sin(x-10\pi)\chi_{[9\pi,11\pi]}(x)\mathbf{r}_{3},$ (72)
and the solution is evolved up to $t_{F}=50s$. Reflecting boundary conditions are set at $x_{\infty}$ by changing the sign of the velocity component of the solution in the rightmost ghost cell. The numerical reflection coefficient is estimated as the ratio
$C^{\text{num}}_{R}=\frac{\max_{x\in[0,x_{s}]}u(t_{F},x)-\min_{x\in[0,x_{s}]}u(t_{F},x)}{\max_{x\in[0,x_{s}]}u(0,x)-\min_{x\in[0,x_{s}]}u(0,x)}.$ (73)
For the far-field operators SDO and S-SDO, we consider sponge layers of lengths $\omega\in\{0.01L,0.1L,L\}$ for $L=2\pi$, and $s_{i}=1$ for all $i\in\{1,2,3\}$. For S-SDO we take $d_{i}$ as in (5), with $\Gamma_{d}=\Gamma_{B}$ (17). We proceed analoglously for SDO, but with $d_{i}\equiv 0$ for $i=1,2$. Different damping rates $\sigma_{i}$ are considered to evaluate their influence on the reflection coefficient. Similarly, for the relaxation methods RM-M and RM, we take the weight function $\Gamma=\Gamma_{B}$ as in (17). Since for these methods the reflection coefficients depend only on the number of cells in the sponge layer, we consider a fixed sponge layer length $\omega=L$ and different refinement levels $\Delta x$. For all the tests, good agreement between the numerical and theoretical reflection coefficients is observed until saturation is attained, as shown in Figures 11 and 11 and discussed below.
While the theoretical estimates could inform the choice of parameters for the ABCs, it is important to note that the conditions under which they were derived are more often than not violated in practice. For instance, in a realistic simulation, additional care is taken to absorb waves at the end of the sponge layer $x_{\infty}$, by imposing the far-field state $\overline{\mathbf{q}}$ in the ghost cells or using constant extrapolation instead of perfectly reflecting boundary conditions. A numerical estimate of the reflection coefficient associated with the use of constant extrapolation (with no sponge layer) is displayed in Figures 11 and 11 (note that different resolutions for the numerical scheme are used in Figure 11, thus constant extrapolation has a slightly different effect for each $\Delta x$). Moreover, waves are not always purely outgoing when they reach the sponge layer, which can be even more impactful in the nonlinear case, where characteristic fields are continuously interacting.
Finally, the theoretical estimates presented here for the SDO and RM techniques are insensitive to the numerical discretization. Most importantly, we derive these estimates under the severe assumption that reflections only occur at the end of the sponge layer $x_{\infty}$, while in practice, reflections continuously take place inside the sponge layer. These discrepancies become evident for the SDO ABC in Figure 11, as the numerical reflection coefficients suddenly saturate as the damping rates $\sigma$ increase, and analogously for the relaxation methods in Figure 11 as the grid is refined. These effects must be taken into account when using the theoretical estimates to guide the choice and optimization of parameters for the ABCs, see e.g. [peric2018analytical, peric2020reducing].
Figure 10: Comparsion of numerical (scattered points) and theoretical (lines) reflection coefficients for the linear equations (48) with SDO (circle markers and dashed lines) and S-SDO ABCs (cross markers and dotted lines) respectively.
Figure 11: Comparsion of numerical (scattered points) and theoretical (lines) reflection coefficients for the linear equations (48) with RM-M (circle markers and dashed lines) and RM ABCs (cross markers and dotted lines) respectively.
Appendix B Construction of damping operators
@@
The construction of the damping operators and the matrix-valued weight function of Sections 2.1, 3.4, and 3.2, relies on the computation of a matrix $\mathbf{M}=\mathbf{R}\bm{\Sigma}\mathbf{R}^{-1}$, where $\bm{\Sigma}=\text{diag}(\sigma_{i})$, and $\mathbf{R}$ contains the right eigenvectors of the flux Jacobian matrix as columns. For the linear equations (48) we have
@@
Appendix C Discretization of integral source term with nonconservative product
@@
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})$ $=\int_{x}^{x_{j}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\sum_{i=j}^{m-1}\int_{x_{i}}^{x_{i+1/2}^{-}}\mathbf{M}\partial_{z}\mathbf{q}\,dz$ (78)
$\quad+\int_{x_{i+1/2}^{+}}^{x_{i+1}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\int_{\{x_{i+1/2}\}}\mathbf{M}\partial_{z}\mathbf{q}\,dz,$ (79)
$=\int_{x}^{x_{j}}\mathbf{M}\sigma_{i}\,dz+\sum_{i=j}^{m-1}\int_{x_{i}}^{x_{i+1/2}^{-}}\mathbf{M}\sigma_{i}\,dz$ (80)
$\quad+\int_{x_{i+1/2}^{+}}^{x_{i+1}}\mathbf{M}\sigma_{i+1}\,dz+\int_{\{x_{i+1/2}\}}\mathbf{M}\partial_{z}\mathbf{q}\,dz,$ (81)
@@
$\int_{\{x_{i+1/2}\}}\mathbf{M}\partial_{z}\mathbf{q}\,dz:=\mu({\{x_{i+1/2}\}})=\int_{0}^{1}\mathbf{M}(\phi(s;\mathbf{q}^{-},\mathbf{q}^{+}))\frac{\partial\phi}{\partial s}(s;\mathbf{q}^{-},\mathbf{q}^{+})ds,$ (82)
where $\phi$ is a Lipschitz continuous function connecting $\mathbf{q}^{-}(t)=\lim_{x\to x_{i+1/2}^{-}}\mathbf{q}(x,t)$ with $\mathbf{q}^{+}(t)=\lim_{x\to x_{i+1/2}^{+}}\mathbf{q}(x,t)$ [le_floch_shock_1989]. Here we choose $\phi$ to be simply a straight line $\phi(s)=s\mathbf{q}^{+}+(1-s)\mathbf{q}^{-}$, and approximate the integral in (82) with the midpoint rule.
Appendix D Entropy stability of the slowing-down operator
One can show that, if only outgoing waves are modified, the slowing-down operator does not destroy the entropy-dissipative nature of linear hyperbolic systems. Indeed, we can consider a general strictly hyperbolic linear system of $N$ PDEs
@@ Appendix D Entropy stability of the slowing-down operator
defined over the real line such that $\lim_{x\to\pm\infty}\mathbf{q}(t,x)=\bf 0$. Let $\mathbf{H}$ be an SPD matrix such that $\mathbf{H}\mathbf{A}$ is symmetric. Then $\mathbf{v}=\frac{1}{2}\mathbf{q}^{T}\mathbf{H}\mathbf{q}$ is a convex entropy function for the system, and $\partial_{t}\int_{\mathbb{R}}\mathbf{v}(\cdot,t)\,dx=0$. Let us analyze the effect of the slowing-down operator applied over the sponge layer $x_{s},x_{\infty}]$. For instance, we slow-down the right-going waves, i.e., we solve
@@
where $\mathbf{S}(x)=\mathbf{R}\bm{\Lambda}_{\mathbf{S}}(x)\mathbf{R}^{-1}$ such that $\mathbf{A}=\mathbf{R}\bm{\Lambda}\mathbf{R}^{-1}$, and the main entries of the diagonal matrix $\bm{\Lambda_{S}}(x)$ are given by
$\tilde{\lambda}_{i}(x)=\begin{cases}\lambda_{i},\quad&\text{if }\lambda_{i}\leq 0,\\ \Gamma_{s}(x)\lambda_{i},\quad&\text{if }\lambda_{i}>0.\end{cases}$ (85)
The scalar smooth slowing-down function $s(x)$ strictly decreases from $1$ to 0 inside the sponge layer $x_{s},x_{\infty}]$ and is constant everywhere else.
Since $\mathbf{A}$ is strictly hyperbolic, one can show that $\mathbf{R}^{T}\mathbf{H}\mathbf{R}$ is diagonal and positive definite. Using this fact, we can prove that $\mathbf{H}\mathbf{S}(x)$ is symmetric and $\mathbf{H}\mathbf{S}^{\prime}(x)$ is symmetric and negative semi-definite for all $x\in\mathbb{R}$, since ${\tilde{\lambda}_{i}}^{\prime}(x)\leq 0$. Therefore, multiplying (84) by $\mathbf{q}^{T}\mathbf{H}$ and integrating in space, we get
$\partial_{t}\int_{\mathbb{R}}\mathbf{v}(\cdot,t)\,dx$ $=-\int_{\mathbb{R}}\mathbf{q}^{T}\mathbf{H}\mathbf{S}(x)\partial_{x}\mathbf{q}\,dx$ (86)
$=\frac{1}{2}\int_{\mathbb{R}}\mathbf{q}^{T}\mathbf{H}\mathbf{S}^{\prime}(x)\mathbf{q}-\partial_{x}\left(\mathbf{q}^{T}\mathbf{H}\mathbf{S}(x)\mathbf{q}\right)\,dx$ (87)
$=\frac{1}{2}\int_{x_{s}}^{x_{\infty}}\mathbf{q}^{T}\mathbf{H}\mathbf{S}^{\prime}(x)\mathbf{q}\,dx\leq 0.$ (88)
@@
[…]
@@
$\frac{1}{2}\partial_{t}\int_{\mathbb{R}}u^{2}\,dx$ $=-\frac{1}{3}\int_{\mathbb{R}}s(x)\partial_{x}u^{3}\,dx=-\frac{1}{3}\int_{\mathbb{R}}\partial_{x}[s(x)u^{3}]-s^{\prime}(x)u^{3}\,dx$ (91)
$=\frac{1}{3}\int_{x_{s}}^{x_{\infty}}s^{\prime}(x)u^{3}\,dx.$ (92)
[…]
While this is not the case for the nonlinear Euler equations (43), where the sign of the wave speeds is independent of the solution, numerical experiments suggest that the slowing-down operator can lead to small amounts of entropy production inside the sponge layer. To illustrate this, we consider the piston problem setup described in Section 5.1. We use $N=250$ cells per wavelength $L$, and a sponge layer of length $\omega=2L$. To isolate the effect of the slowing-down operator, we run a simulation with both slowing-down and damping operators, and one with only the slowing-down operator only. In Figure 12 we can see that, in the absence of a damping operator, slowing-down leads to wave compression inside the sponge layer. While the shock-capturing nature of the scheme is able to stabilize the solution, large spurious oscillations propagate back into the domain. In Figure 13 we can see the evolution of the entropy function (see e.g. [whitham_linear_1974, Sec. 6.4])
$s(\mathbf{q})=\frac{1}{1-\gamma}\log\left(pV^{\gamma}\right),$ (93)
integrated over the sponge layer $x_{s},x_{\infty}]$. When only the slowing-down operator is used, there are periodic increments of entropy inside the sponge layer, which are controlled by the numerical viscosity of the discretization. However, as numerical diffusion vanishes with mesh refinement or the use of higher-order schemes, this behavior could lead to overall instability.
Figure 12: Effect of the slowing-down operator for the nonlinear system (43).
Figure 13: Time evolution of the entropy function (93) integrated over the sponge layer $x_{s},x_{\infty}]$